Ch 2-3 Exercises& Problems Account Balances a. During February, $96,940 was paid to creditors on account, and purchases on account were $124,080. Assuming that the February 28 balance of Accounts Payable was $41,680, determine the account balance on February 1. b. On October 1, the accounts receivable account balance was $48,000. During October, $417,600 was collected from customers on account. Assuming that the October 31 balance was $55,200, determine the fees billed to customers on account during October. c. On April 1, the cash account balance was $18,120. During April, cash receipts totaled $275,420 and the April 30 balance was $13,050. Determine the cash payments made during April. 6:49 PM 124/2019
Answer #1

1.Beginning Balance(Feb1) + Purchases – Payments on Purchases = Ending Balance

Beginning Balance + 124,080 -96,940 = 41680

Beginning Balance = 14,540

2. Beginning Balance +Credit sales – Collections = Ending Balance

48000+ credit sales – 417,600 = 55200

Credit sales = 424,800

3. Beginning Balance + Cash receipts – Cash payments = Ending Balance

18,180 +275,420 – Cash Payment = 13,050

= 293,600 – Cash Payment = 13050

= 293600 – 13050 = 280550

Cash Payment = 280,550